# Neural Thermodynamic Integration: Free Energies from Energy-based Diffusion Models

### v2

@@ Abstract
Thermodynamic integration (TI) offers a rigorous method for estimating free-energy differences by integrating over a sequence of interpolating conformational ensembles. However, TI calculations are computationally expensive and typically limited to coupling a small number of degrees of freedom due to the need to sample numerous intermediate ensembles with sufficient conformational-space overlap. In this work, we propose to perform TI along an alchemical pathway represented by a trainable neural network, which we term Neural TI. Critically, we parametrize a time-dependent Hamiltonian interpolating between the interacting and non-interacting systems, and optimize its gradient using a denoising-diffusion objective. The ability of the resulting energy-based diffusion model to sample all intermediate ensembles allows us to perform TI from a single reference calculation. We apply our method to Lennard-Jones fluids, where we report accurate calculations of the excess chemical potential, demonstrating that Neural TI is capable of coupling hundreds of degrees of freedom at once.
@@
Figure 1: Schematic summary of the proposed approach. We propose to interpolate between the target, $\mathcal{H}_{0}$, and latent, $\mathcal{H}_{1}$, Hamiltonians with a time-dependent potential $U_{t}^{\theta}$. During sampling, the normalizing constant of the target can be estimated via thermodynamic integration. Particles whose separation from their closest neighbor is less than $0.85\sigma$ are colored red, the rest are colored blue. This color-coding illustrates that in the ideal gas (right) there are many colliding particles and as the LJ potential is turned on (left) the particles do not overlap anymore.
@@ I Introduction
Machine learning—in particular generative models—is rapidly transforming how we model molecular systems [6, 7, 8, 9]. Recently, normalizing flows have been proposed for sampling in the context of statistical physics [10], quantum field theory [11], and molecules [12]. Flow-based approaches are attractive as they represent exact probability densities that can be used for unbiased estimation of observables when combined with importance sampling to correct for the mismatch between the learnt and target densities. Flow-based methods have been proposed as a means to compute free-energy differences [12, 13]. However, flows also present challenges: coupling flows [14] are cumbersome to work with when one needs to encode physical bias in the architecture, and continuous flows [15] suffer from the cost of computing the divergence of the model.
In this work, we instead propose to use denoising diffusion models (DDMs) [16, 17] to estimate free-energy differences. To this end, we assign our data and latent distributions to the interacting and non-interacting Hamiltonians, respectively. We then train a time-dependent, interpolating potential $U_{t}^{\theta}$ between the potentials of the target, $U_{0}$, and the prior, $U_{1}$, along the finite-time interval of the DDM. We match the force exerted by the potential to the Hyvärinen score of the DDM
@@
We demonstrate our methodology on computer simulations of a condensed-phase, many-body system: the Lennard-Jones (LJ) liquid. The particles are confined to a box with periodic boundary conditions, leading to topological constraints on the DDM addressed further below. The latent space consists of the ideal-gas system: particles that carry kinetic energy, but do not interact. The DDM thereby interpolates between interacting and non-interacting system, whose trajectories model the dynamics of a time-dependent force-field $-\nabla U_{t}^{\theta}$. See Figure 1 for a schematic summary of the proposed approach. We validate Neural TI on accurate calculations of the excess chemical potential. For the first time, we report accurate free-energy differences of coupling an entire liquid, totalling 600 degrees of freedom.
II Thermodynamic integration
@@
where $\langle\partial_{\lambda}U_{\lambda}\rangle_{\lambda}$ denotes the expected value of $\partial_{\lambda}U_{\lambda}$ under the density $\rho_{\lambda}=Z_{\lambda}^{-1}\text{e}^{-\beta U_{\lambda}}$. Practically, the small phase-space overlap between ensembles requires an interpolation of many intermediate Hamiltonians parametrized by the coupling variable $\lambda$ [5]. Here instead we will use a DDM (§III) to learn the alchemical pathway, i.e., we replace $\lambda$ by the diffusion model’s time variable, $t$. The generative model’s ability to sample all intermediate ensembles will do away with the need for intermediate reference simulations, and provide an accurate estimate of much larger free-energy differences than reported so far [18].
III Denoising Diffusion models
Diffusion models [16, 17] are a class of generative models defined by a pair (forward and reverse) of stochastic processes. In the continuous-time formulation [19], the forward process is given by the SDE
@@ III Denoising Diffusion models
where $\text{d}W_{t^{\rightarrow}}$ is a Wiener process and SDE starts from the initial condition $X_{0}\sim\rho_{0}$. Note that $\rho_{t}(x_{t}|x_{0})$ is a Gaussian for all $t$ and $x_{0}$. In particular $f_{t}$ and $g_{t}$ are time-dependent and are chosen such that $\rho_{t}(x_{1}|x_{0})$ is close to a standard Gaussian distribution for all $x_{0}$. The noising process then corresponds to simulating Eq. (6) from $t=0$ to $t=1$. The same time marginals can be reproduced by a time-reversed SDE [20] from $t=1$ to $t=0$,
@@
with initial condition $Y_{0}\sim\mathcal{N}(0,\mathbb{I})$. Given the score $s(x,t)=\nabla_{x}\log\rho_{t}(x)$, one could use the reverse Eq. (7) to map Gaussian samples to the initial target density $\rho_{0}$.
@@
III.1 TI with diffusion models
@@
III.2 Diffusion models on tori
In what follows we will be interested in particles living in a $d$-dimensional box with periodic boundary conditions. To work with such systems the usual framework of diffusion models needs to slightly adjusted to accommodate for the different topology of the configurational space. Topologically speaking, the position of a single particle is specified by a point on a hypertorus, $\mathbb{T}^{N}$, of dimension $d$, and the configurational space of $N$ particles is then a $dN$-dimensional hypertorus, $\mathbb{T}^{dN}$.
@@ III.2 Diffusion models on tori
Jing et al. [25] compute the score of the wrapped Gaussian in Eq. (12) by truncating both the numerator and denominator of the logarithmic derivative $\nabla\log\rho_{t}(x_{t}|x_{0})=\frac{\nabla\rho_{t}(x_{t}|x_{0})}{\rho_{t}(x_{t}|x_{0})}$. Alternatively, one could write the score at $x_{t}$ of the wrapped Gaussian as a weighted average of the scores of the unwrapped Gaussian over the fiber above $x_{t}$
@@ IV.1 Canonical ensemble ($NVT$)
In the ideal gas particles do not interact with each other and the system is described by a Hamiltonian only containing a kinetic term $\mathcal{H}_{\text{ideal}}(\mathbf{p})=\sum_{i}\frac{p_{i}^{2}}{2m}$. In this case $Z_{N}$ can be analytically computed
@@
where $V$ is the volume of the box and $\Lambda=h/\sqrt{2\pi m\beta^{-1}}$ is the thermal wavelength. Since particle positions in the ideal gas are independently and uniformly distributed over the box, the latent space of the toroidal diffusion model (§III.2) describes the positions $\mathbf{q}$ of an ideal gas in the canonical ensemble.
@@
IV.2 Grand Canonical ensemble ($\mu VT$)
@@
Intuitively, the excess chemical potential $\mu_{\text{ex}}=\mu-\mu_{\text{ideal}}$ measures the deviation in chemical potential due to the interaction term of the Hamiltonian at fixed density.
The straightforward way of estimating the excess chemical potential of an interacting system is to perform a grand-canonical Monte Carlo (GCMC) simulation on the distribution in Eq. (22) at a prescribed value of $\mu$. Grand-canonical sampling of the number of particles allows us to estimate the empirical mean $\langle N\rangle_{\text{GCMC}}$, from which we extract $\mu_{\text{ideal}}$ via Eq. (26). From $\mu$ and $\mu_{\text{ideal}}$, one can simply compute the excess chemical potential, $\mu_{\text{ex}}=\mu-\mu_{\text{ideal}}$.
In this work we estimate the excess chemical potential from several canonical simulations. We train a generative model on several canonical ensembles and then estimate the canonical partition function $Z_{N}$ for each $N$. The collection $\{Z_{N}\}$ is used to compute the distribution of the number of particles, $p(N)$, see Eq. (24), for a predefined choice of the chemical potential, $\mu$. In combination with Eq. (26), we compute the excess chemical potential, $\mu_{\text{ex}}$.
V Application: Lennard-Jones fluid
We consider a crowded condensed-matter system: a collection of three-dimensional particles confined in a box, interacting via a pairwise Lennard-Jones (LJ) potential
@@ V Application: Lennard-Jones fluid
    Estimate the partition function, $Z_{N}$, by generating configurations at a given $N$ from the DDM and computing $\hat{Z}_{N}$ from TI, see Eq. (21);
@@
where $U_{t}^{\text{LJ}}$ is a soft-core LJ-potential [27] with a time-dependent softening parameter
@@
This soft-core potential ensures that the numerically unstable region of the LJ-potential is not evaluated on noisy samples but is slowly introduced as more and more noise is removed from the samples (Figure 2). The only trainable component, $U^{\theta}_{t}$ can then be parametrized using ideas from the force-field literature [28, 29, 30, 31]. Strictly speaking, our model learns a time-dependent potential, and thus the above architectures are only applicable once conditioned on the diffusion time, $t$.
Figure 2: The soft-core LJ potential $U_{t}^{\text{LJ}}$ in Eq. (29) for various values of $t\in[0,1]$. Note that for larger values of $t$, particles can get closer to each other without experiencing strong repulsive forces. This is necessary since in the diffusion process it is inevitable that particles get close to each other as $t$ increases.
@@ Results
We set the dimension of the system to $d=3$, the volume of the box to $V=216$, the inverse temperature to $\beta=1$, the mass of the particles to $m=1$, Planck’s constant to $h=10^{-4}$, and the parameters of the LJ-potential to $\varepsilon=0.8,\sigma=1$. We train a single DDM on samples from canonical Monte Carlo simulations at particle numbers $N\in\{40,80,120,160,200\}$ (i.e. densities $\rho=N/V\in\{0.19,0.37,0.56,0.74,0.93\}$). We refer the reader to Appendix B for details on the Monte Carlo simulations and on the architecture.
To evaluate the generative performance of these models we compare the radial distribution function (RDF), $g(r)$, between Monte Carlo samples and samples from the trained DDM (Figure 3). We find accurate reconstructions across a broad range of densities, $\rho=0.19$ to $\rho=0.93$. The shapes of the RDF clearly indicate a transition from gas to liquid as we increase particle density.
Figure 3: Radial distribution functions as predicted by Monte Carlo simulations (gray) and a diffusion model (red) trained on densities $\rho\in\{0.19,0.37,0.56,0.74,0.93\}$. Note that the model reconstructs $g(r)$ across the the gas-liquid phase transition. A Gaussian kernel with $\sigma=0.03$ was applied to all curves in this plot.
@@
Figure 4: Distribution of the number of particles in the grand canonical ensemble at different chemical potentials from GCMC simulations (gray) and estimated by thermodynamic integration with a trained diffusion model (red). The vertical blue lines denote the canonical ensembles that the diffusion model was trained on. A Gaussian kernel with $\sigma=0.015$ was applied to all curves in this plot.
Further, we also evaluate the functional relationship between $\rho$ and $\mu$ in Figure 5. The left subfigure shows the average density as a function of the chemical potential, as calculated by both reference GCMC and our proposed TI methodology. Though the reference calculations are run in the grand-canonical ensemble, we make use of TI from a set of canonical simulations to estimate $\mu$. The TI calculations estimate the partition functions, $\hat{Z}_{N}$, at different values of $N$. From those, we compute the particle-number distribution via Eq.(24), extract the mean, and deduce the average particle density, $\langle\rho\rangle$. We find excellent agreement across the gas–liquid phase transition—visually illustrated by the jump in density. The right subfigure shows the excess chemical potential as a function of particle density. Our methodology allows us to smoothly interpolate across the phase transition. At high density, the slight underestimation of the particle-number distribution leads a small discrepancy in the excess chemical potential. The main workhorse-method to compute excess chemical potentials from canonical simulations, the Widom particle-insertion method, shows significant difficulties in this regime due to its perturbative nature [32, 33].
@@
Neural TI offers unique insight in the equilibrium properties of liquids by directly reporting the free energy of coupling the entire liquid. Figure 6 shows the coupling free energy from the ideal gas to the target Hamiltonian of the LJ particle configurations at different densities. On the left-most part, values close to 0 density lead to virtually no free energy due to the ideal-gas limit. As density increases close to $\rho\approx 0.75$, we observe a maximum—interestingly quite a bit higher in density compared to where the excess chemical potential attains its minimum, i.e., $\langle\rho\rangle\approx 0.55$. Instead, we hypothesize that the peak occurs where significant particle-overlap starts severely hampering the configurational space available—this is exactly the regime where perturbative methods, e.g., Widom insertion, become challenging. At a density of $1.0$, the available configurational integral reduces significantly—as can be seen by the drop to low $\Delta\hat{F}$ values. We finally point at the sheer amplitude of the estimated free-energy differences: up to $200\leavevmode\nobreak\ k_{\text{B}}T$ in free energies are reported. Traditional TI methodologies require tens of interpolating Hamiltonians, such that each reference simulation contributes on the order of $k_{\text{B}}T$ each. Neural TI’s ability to reach two orders of magnitude more demonstrates the appeal to learn the alchemical pathway. While traditional alchemical-transformation methods focus on minimizing the change in Hamiltonian to a handful of degrees of freedom, here we report accurate free-energy differences of coupling an entire box of LJ particles, totaling 600 degrees of freedom.
@@
Figure 6: Neural TI-estimated free energy cost of turning on the LJ interactions in the canonical ensemble at different densities. The subscript $1\rightarrow 0$ corresponds to an interpolation from the non-interacting latent space to the target Hamiltonian, i.e., turning on interactions in the system.
VI Conclusion

### v3

@@ Abstract
Thermodynamic integration (TI) offers a rigorous method for estimating free-energy differences by integrating over a sequence of interpolating conformational ensembles. However, TI calculations are computationally expensive and typically limited to coupling a small number of degrees of freedom due to the need to sample numerous intermediate ensembles with sufficient conformational-space overlap. In this work, we propose to perform TI along an alchemical pathway represented by a trainable neural network, which we term Neural TI. Critically, we parametrize a time-dependent Hamiltonian interpolating between the interacting and non-interacting systems, and optimize its gradient using a score matching objective. The ability of the resulting energy-based diffusion model to sample all intermediate ensembles allows us to perform TI from a single reference calculation. We apply our method to Lennard-Jones fluids, where we report accurate calculations of the excess chemical potential, demonstrating that Neural TI reproduces the underlying changes in free energy without the need for simulations at interpolating Hamiltonians.
@@
Figure 1: Schematic summary of the proposed approach. We interpolate between the target, $\mathcal{H}_{0}$, and latent, $\mathcal{H}_{1}$, Hamiltonians with a time-dependent potential $U_{t}^{\theta}$. During sampling, the normalizing constant of the target can be estimated via thermodynamic integration. Particles whose separation from their closest neighbor is less than $0.85\sigma$ are colored red, the rest are colored blue. This color-coding illustrates that in the ideal gas (right) there are many colliding particles and as the LJ potential is turned on (left) the particles do not overlap anymore.
@@ I Introduction
Machine learning—in particular generative models—is rapidly transforming how we model molecular systems [6, 7, 8]. Normalizing flows have been proposed for sampling in the context of statistical physics [9], lattice quantum field theory [10], and molecules [11]. Flow-based approaches are attractive as they represent exact probability densities that can be used for unbiased estimation of observables when combined with importance sampling to correct for the mismatch between the learnt and target densities. Invernizzi et al. [12] have revisited parallel tempering by a learned flow map to increase phase-space overlap. Flow-based approaches have also been extended to estimate free energy differences via ML-enhanced free energy perturbation [13] and generalized to different thermodynamical ensembles [14]. Flows, however, also present challenges: coupling flows [15] are cumbersome to work with when one needs to encode physical bias in the architecture, and continuous flows [16] suffer from the cost of divergence computations.
In this work, we instead propose to use denoising diffusion models (DDMs) [17, 18] to estimate free-energy differences. To this end, we assign our data and latent distributions to the interacting and non-interacting Hamiltonians, respectively. We then train a time-dependent, interpolating potential $U_{t}^{\theta}$ between the potentials of the target, $U_{0}$, and the prior, $U_{1}$, along the finite-time interval of the DDM. We match the force exerted by the potential to the (Stein) score [19] of the DDM
@@
We demonstrate our methodology on computer simulations of a condensed-phase, many-body system: the Lennard-Jones (LJ) liquid. The particles are confined to a box with periodic boundary conditions, leading to topological constraints on the DDM addressed further below. The latent space consists of the ideal-gas system: particles that carry kinetic energy, but do not interact. The DDM thereby interpolates between interacting and non-interacting system, whose interpolating densities correspond to the Boltzmann densities of time-dependent machine learned potential $U_{t}^{\theta}$. See Figure 1 for a schematic summary of the proposed approach. We validate Neural TI on accurate calculations of the excess chemical potential. We report accurate free-energy differences of coupling all LJ interactions starting from the ideal gas, without relying on simulations at interpolating Hamiltonians.
Thermodynamic integration
@@
$=-\int_{0}^{1}\text{d}\lambda\,\frac{1}{Z_{\lambda}}\partial_{\lambda}Z_{\lambda}$ (3)
@@
where $\langle\partial_{\lambda}U_{\lambda}\rangle_{\lambda}$ denotes the expected value of $\partial_{\lambda}U_{\lambda}$ under the density $\rho_{\lambda}=Z_{\lambda}^{-1}\text{e}^{-\beta U_{\lambda}}$. Practically, the small phase-space overlap between ensembles requires an interpolation of many intermediate Hamiltonians parametrized by the coupling variable $\lambda$ [5]. Here instead we will use a DDM to learn the alchemical pathway, i.e., we replace $\lambda$ by the diffusion model’s time variable, $t$. Diffusion models allow us to sample the equilibrium distribution of each intermediate state, meaning that our approach does not coincide with non-equilibrium variants, such as slow- or fast-growth TI [20]. The generative model’s ability to sample all intermediate ensembles will do away with the need for intermediate reference simulations, and provide an accurate estimate of much larger free-energy differences than reported so far [21]. In the machine-learning community, Masrani et al. [22] used TI for tightening the evidence lower bound (ELBO).
Denoising Diffusion models
Diffusion models [17, 18] are a class of generative models defined by a pair (forward and reverse) of stochastic processes. In the continuous-time formulation [23], the forward process is given by the stochastic differential equation (SDE)
@@
where $\text{d}W_{t^{\rightarrow}}$ is a Wiener process and SDE starts from the initial condition $X_{0}\sim\rho_{0}$. Note that $\rho_{t}(x_{t}|x_{0})$ is a Gaussian for all $t$ and $x_{0}$. In particular $f_{t}$ and $g_{t}$ are time-dependent and are chosen such that $\rho_{t}(x_{1}|x_{0})$ is close to a standard Gaussian distribution for all $x_{0}$. The noising process then corresponds to simulating Equation 7 from $t=0$ to $t=1$. The same time marginals can be reproduced by a time-reversed SDE [24] from $t=1$ to $t=0$,
@@
with initial condition $Y_{0}\sim\mathcal{N}(0,\mathbb{I})$. Given the score $s(x,t)=\nabla_{x}\log\rho_{t}(x)$, one could use the reverse Equation 8 to map Gaussian samples to the initial target density $\rho_{0}$.
@@
TI with diffusion models
@@
Diffusion models on tori
In what follows we will be interested in particles living in a $d$-dimensional box with periodic boundary conditions. To work with such systems the usual framework of diffusion models needs to be slightly adjusted to accommodate for the different topology of the configurational space. Topologically speaking, the position of a single particle is specified by a point on a hypertorus, $\mathbb{T}^{N}$, of dimension $d$, and the configurational space of $N$ particles is then a $dN$-dimensional hypertorus, $\mathbb{T}^{dN}$.
@@
[…]
@@ II Statistical ensembles
In the ideal gas particles do not interact with each other and the system is described by a Hamiltonian only containing a kinetic term $\mathcal{H}_{\text{ideal}}(\mathbf{p})=\sum_{i}\frac{p_{i}^{2}}{2m}$, where $m=m_{1}=...=m_{N}$ denotes the mass of all particles. In this case $Z_{N}$ can be analytically computed
@@
where $V$ is the volume of the box and $\Lambda=h/\sqrt{2\pi m\beta^{-1}}$ is the thermal wavelength. Since particle positions in the ideal gas are independently and uniformly distributed over the box, the latent space of the toroidal diffusion model describes the positions $\mathbf{q}$ of an ideal gas in the canonical ensemble.
@@
Grand Canonical ensemble ($\mu VT$)
@@
Intuitively, the excess chemical potential $\mu_{\text{ex}}=\mu-\mu_{\text{ideal}}$ measures the deviation in chemical potential due to the interaction term of the Hamiltonian at fixed density. The straightforward way of estimating the excess chemical potential of an interacting system is to perform a grand-canonical Monte Carlo (GCMC) simulation on the distribution in Equation 23 at a prescribed value of $\mu$. Grand-canonical sampling of the number of particles allows us to estimate the empirical mean $\langle N\rangle_{\text{GCMC}}$, from which we extract $\mu_{\text{ideal}}$ via Equation 27. From $\mu$ and $\mu_{\text{ideal}}$, one can simply compute the excess chemical potential, $\mu_{\text{ex}}=\mu-\mu_{\text{ideal}}$.
In this work we estimate the excess chemical potential from several canonical simulations. We train a generative model on several canonical ensembles and then estimate the canonical partition function $Z_{N}$ for each $N$. The collection $\{Z_{N}\}$ is used to compute the distribution of the number of particles, $p(N)$, see Equation 25, for a predefined choice of the chemical potential, $\mu$. In combination with Equation 27, we compute the excess chemical potential, $\mu_{\text{ex}}$.
III Application: Lennard-Jones fluid
We now consider a many-body condensed-phase system: a collection of three-dimensional particles confined in a box, interacting via a pairwise Lennard-Jones (LJ) potential
@@ III Application: Lennard-Jones fluid
    Estimate the partition function, $Z_{N}$, by generating configurations at a given $N$ from the DDM and computing $\hat{Z}_{N}$ from TI, see Equation 22;
@@
where $U_{t}^{\text{LJ}}$ is a soft-core LJ potential [31] with a time-dependent softening parameter
@@
This soft-core potential ensures that the numerically unstable region of the LJ potential is not evaluated on noisy samples but is slowly introduced as more and more noise is removed from the samples (Figure 2). The only trainable component, $U^{\theta}_{t}$ can then be parametrized using ideas from the machine learning force field literature [32, 33, 34, 35]. The parametrization in Equation (29) means that the neural network learns the deviation from the pathway typically used in the standard approach to TI.
Figure 2: The soft-core LJ potential $U_{t}^{\text{LJ}}$ in Equation 30 for various values of $t\in[0,1]$. Note that for larger values of $t$, particles can get closer to each other without experiencing strong repulsive forces. This is necessary since in the diffusion process it is inevitable that particles get close to each other as $t$ increases.
@@ Results
We set the dimension of the system to $d=3$, the volume of the box to $V=216$, the inverse temperature to $\beta=1$, the mass of the particles to $m=1$ and the parameters of the LJ potential to $\varepsilon=0.8,\sigma=1$. We set the value of Planck’s constant to $10^{-4}$, but stress that the observables we report do not depend on $h$, its numerical value was chosen out of convenience. We train a single DDM on samples from canonical Monte Carlo simulations at particle numbers $N\in\{40,80,120,160,200\}$ (i.e. densities $\rho=N/V\in\{0.19,0.37,0.56,0.74,0.93\}$). We refer the reader to Appendix B for details on the Monte Carlo simulations and on the architecture.
To evaluate the generative performance of these models we compare the radial distribution function (RDF), $g(r)$, between Monte Carlo samples and samples from the trained DDM (Figure 3). We find accurate reconstructions across a broad range of densities, $\rho=0.19$ to $\rho=0.93$. The shapes of the RDF clearly indicate a transition from gas to liquid as we increase particle density. Note that reconstructing observables from generated samples hints at a good fit between the learnt and target densities. It is a necessary but not sufficient condition for the accuracy of TI along the model.
Figure 3: Radial distribution functions as predicted by Monte Carlo simulations (gray) and a diffusion model (red) trained on densities $\rho\in\{0.19,0.37,0.56,0.74,0.93\}$. Note that the model reconstructs $g(r)$ across the the gas-liquid phase transition.
@@
Figure 4: Distribution of the number of particles in the grand canonical ensemble at different chemical potentials from GCMC simulations (gray) and estimated by thermodynamic integration with a trained diffusion model (red). The vertical blue lines denote the canonical ensembles that the diffusion model was trained on.
Further, we also evaluate the functional relationship between $\rho$ and $\mu$ in Figure 5. The left subfigure shows the average density as a function of the chemical potential, as calculated by both reference GCMC and our proposed TI methodology. Though the reference calculations are run in the grand-canonical ensemble, we make use of TI from a set of canonical simulations to estimate $\mu$. The TI calculations estimate the partition functions, $\hat{Z}_{N}$, at different values of $N$. From those, we compute the particle-number distribution via Equation 25, extract the mean, and deduce the average particle density, $\langle\rho\rangle$. We find excellent agreement across the gas–liquid phase transition—visually illustrated by the jump in density. The right subfigure shows the excess chemical potential as a function of particle density. Our methodology allows us to smoothly interpolate across the phase transition. At high density, the slight underestimation of the particle-number distribution leads a small discrepancy in the excess chemical potential. The main workhorse-method to compute excess chemical potentials from canonical simulations, the Widom particle-insertion method, shows significant difficulties in this regime due to its perturbative nature [36, 37].
@@
This experiment also offers insight in the equilibrium properties of liquids by directly reporting the free energy of coupling all LJ interactions starting from the ideal gas. Figure 6 shows the coupling free energy starting from the ideal gas to the target Hamiltonian of the LJ particle configurations at different densities. On the left, densities close to 0 lead to virtually no free energy due to the ideal-gas limit. As density increases close to $\rho\approx 0.75$, we observe a maximum—interestingly quite a bit higher in density compared to where the excess chemical potential attains its minimum, i.e., $\langle\rho\rangle\approx 0.55$. Instead, we hypothesize that the peak occurs where significant particle-overlap starts severely hampering the configurational space available—this is exactly the regime where perturbative methods, e.g., Widom insertion, become challenging. At a density of $1.0$, the available configurational integral reduces significantly—as can be seen by the drop to low $\Delta\hat{F}$ values.
For comparison, we perform the same estimate using standard TI with $5,10,20,50$ and $100$ evenly spaced intermediate simulations. The convergence of the standard TI curves with the number of intermediate simulations validates Neural TI as a free-energy estimator. At high densities the slight discrepancy between Neural TI and standard TI with many simulations is likely both due to convergence issues of the reference simulations and limited expressivity of our time-dependent ML force-field architecture.
We finally point at the sheer amplitude of the estimated free-energy differences: up to $200\leavevmode\nobreak\ k_{\text{B}}T$ in free energies are reported. Traditional TI methodologies require tens of interpolating Hamiltonians. Neural TI’s ability to estimate free energy differences from a single reference simulation demonstrates the appeal to learn the alchemical pathway. While traditional alchemical-transformation methods focus on minimizing the change in Hamiltonian to a handful of degrees of freedom, here we report free-energy differences of coupling all LJ interactions to a box of ideal-gas particles.
@@
Figure 6: Estimated free energy of turning on the LJ interactions in the canonical ensemble at different densities (number of particles at the top). The subscript $1\rightarrow 0$ corresponds to an interpolation from the non-interacting latent space to the target Hamiltonian, i.e., turning on interactions in the system. Estimate by Neural TI shown in thick black. For comparison, we also display various standard TI estimates with increasing numbers of reference simulations (various colors).
IV Discussion and Limitations
Possibly the simplest way to summarize the idea behind this work is to state that energy-based parametrization combined with TI extracts a likelihood estimate from diffusion models at a relatively low cost. Given the convergence of continuous-time normalizing flows and diffusion models [10, 38], it is worth comparing our approach to Boltzmann generators [11]. The key difference is that, unlike our approach, flows provide an exact likelihood of the model density that can later be used for importance sampling. This exact likelihood, however, requires the divergence of a vector field, which is considerably more expensive to compute than the temporal derivative of a scalar energy. Diffusion models equipped with neural TI thus offer an appealing alternative to flow-based sampling of molecular systems: an approximate likelihood at low cost.
We can also think of the proposed approach as thermodynamic integration with an arbitrary number of intermediate $\lambda$-slices, inversely proportional to the step size of the integrator of the reverse dynamics. As with standard TI, using more steps increases accuracy but also computational cost. Our approach, however, bypasses the burden of setting up intermediate simulations, as the reverse dynamics can generate i.i.d. samples from all intermediate distributions once the model is trained. The computational tradeoff is thus between performing the intermediate simulations and training a diffusion model. In our experiment on the LJ system, neither neural TI nor standard TI was optimized for speed, and required comparable times to perform on the same hardware. We leave a detailed analysis of the computational aspects and scaling properties of the method for future work.
Looking towards chemically relevant applications, molecular systems contain a range of important physico-chemical interactions. Here we demonstrated the applicability of Neural TI to LJ interactions—a prominent model for repulsion and dispersion interactions. Extensions of the method to intramolecular interactions, as well as long-range Coulomb electrostatics, would require an adaptation of the energy function in Equation 29.
Conclusion